OpenStudy (anonymous):

A particle moves along a straight line with equation of motion s = f(t), where s is measured in meters and t in seconds. Find the velocity and speed when t = 4. f(t) = 50 + 40t − 4.5t^2

OpenStudy (ash2326):

Sorry for that guys, some troll is on rampant.

OpenStudy (anonymous):

Ok, just f(t) = 50 + 40t − 4.5t^2 the first derivate so, f'(t)=−9t+ 40 f'(t)=−9(4)+ 40 f'(t)=−36+ 40 f'(t)= 4 If you need the aceleration is the second derivate f''(t)=−9

OpenStudy (anonymous):

If you're allowed to use calculus: Velocity is the derivative of displacement: v(t) = x'(t) f'(t) = 40-9t f'(4) = 40 - 9(4) = 40 - 36 = 4 and speed is just the absolute value of velocity, so also four.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Erick was answering the question when I started to, so I feel he deserves a medal too. so I'm gonna give you a best answer :)

OpenStudy (anonymous):

@misha511 or @Erick.04, would I be right in saying to get the derivative I have to use f((x+h) - f(x)) / h? I'm a little lost on setting up that derivative. And I wish they let you give out more than one medal :c

OpenStudy (anonymous):

Ah! Figured it out! Thanks guys